The resistance of a platinum resistance thermometer is 10 ohms at 0oC and 10.4 ohms at 1000C. What is the temperature when the resistance of the platinum is 10.30 ohms?
1
Expert's answer
2020-05-28T13:31:20-0400

The equation for the resistance of a platinum resistance is linear:


"R(T) = kT + b"

where "T" is the temperature and "k, b" are coefficients. According to the provided data:


"R(0) = k\cdot 0 + b = 10\Omega \Rightarrow b = 10\Omega"

and:


"R(1000\degree C) = k\cdot 1000\degree C + 10\Omega = 10.4\Omega \Rightarrow k = 0.4\cdot 10^{-3}\Omega/\degree C"

Thus, when the resistance is 0.30 ohms, the temperature will be:


"R(T) = 0.4\cdot 10^{-3}\cdot T + 10\Omega = 10.3\Omega \\
T = 750 \degree C"

Answer. T = 750 C.